OpenStudy (anonymous):

I need help determing my final grade. Homework/quizzes= 25% of final grade Final Exam= 25% of final grade Tests 1,2,3 each count 14.3% and the Mastering Exam counts 7.1%, and they add up to 50% of the Final Grade. I have 558.57/637= 87.69% on the homework/quizzes. I have an 85% on the Mastering Exam. I have 71.09% on the three tests combined. What is my grade without the Final Exam in?

OpenStudy (anonymous):

if 637 pts is 25% of your grade the total amount of points you can get is \[637*4 =2548\] each test counts 14.3% \[.143*2548*3=1093.1 pts \] 71.09% of that will get your points from you tests \[.7109*1093.1=777.1****\]

OpenStudy (anonymous):

7.1% comes from mastering examp \[.071*2548=180.91\] if you got 85% on that \[.85*180.91=153.8*******\]

OpenStudy (anonymous):

so then you have test points +masterin exam points + homework quizzes \[777.1+153.8+558.57=1489.47\] \[\frac{1489.47}{2548}=58% \] this is ofcourse including your final do you want me to not include it for your grade at this moment?

OpenStudy (anonymous):

not including the final.. I won't take that until next Thursday. I want to know my current grade and what grade I need on my Final to receive a B.

OpenStudy (anonymous):

\[2548-637=1911\] \[\frac{1489.47}{1911}=78 percent\]

OpenStudy (anonymous):

right now you have a 78 %

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

you'd need a 86% to pass with a b

OpenStudy (anonymous):

UGH THAT'S HORRIBLE ;( thanks...

OpenStudy (anonymous):

yeah if you didn't take the final with a 100% you'd get a solid c... the final exam is weighted pretty high =/

OpenStudy (anonymous):

and the tests were about half your grade

OpenStudy (anonymous):

Yeah, the class is Calculus for Engineers. My professor decided to tell us today that the grades that were in our database aren't our actual grades... On there I have an 83%...

OpenStudy (anonymous):

if you had averaged like a 75% in tests you'd not need as much =/

OpenStudy (anonymous):

I know.. I have horrible test anxiety and needing an 85% on my final is gonna be hell for me.

OpenStudy (anonymous):

did you get a test back recently? that's probably what dropped it so much =/

OpenStudy (anonymous):

Yeah the third test I got it back Yesterday.. I got 70/100

OpenStudy (anonymous):

if you need help studying calculus, you can ask me @jim_thompson5910 , @UnkleRhaukus that's lame he grades 70/100 and doesn't give you the actual points per each so it's easier on you to find your grade =/

OpenStudy (anonymous):

I still have homework as well to go in.. 77 more points into the 25% category and I usually do well on the homework.

OpenStudy (anonymous):

oh well then that'll help a ton haha let me check if i subtract 77 out of the final

OpenStudy (anonymous):

subtract 77?

OpenStudy (anonymous):

can you figure out what my grade is if I get 100% on the entire 77 points that still needs to be added into the homework category?

OpenStudy (anonymous):

74% =]

OpenStudy (anonymous):

74% what?

OpenStudy (anonymous):

o wait nvm let me do this again lol

OpenStudy (anonymous):

hahaha! (;

OpenStudy (anonymous):

if you do all your homework and get 77points, you'll end up with an 82% before the final exam then you'll need an 75% on the final to end up with a 80%

OpenStudy (anonymous):

77 pts is almost like 5% of your grade hah so it makes a big difference

OpenStudy (anonymous):

phew. Okay

OpenStudy (anonymous):

That gives me a little more room to f**k up haha... I have 4 exams that day and I am stressing big time. I never get C's... So to get a C..... ----> I die. Hahah.

OpenStudy (anonymous):

Plus, I'm on a scholarship and I have to receive to certain gpa to keep it. ;/ ohhhh life of an engineering major..... haha

OpenStudy (anonymous):

<---- Dual majoring in Computer/Electrical engineering =]

OpenStudy (anonymous):

Ohh very nice! (: I'm majoring in Aerospace Engineering specifically Astronautics. But, recently I've been considering nursing school after I receive my BSE.

OpenStudy (anonymous):

Aerospace is pretty cool, teacher was telling about blackbirds yesterday (eng physics)

OpenStudy (anonymous):

nice! Aerospace is the best! (;

OpenStudy (anonymous):

you probably have no idea what i'm taling about haha

OpenStudy (anonymous):

isn't a blackbird a vehicle that is used in Physics?

OpenStudy (anonymous):

Idk if you're a gamer , but if you've played Call of Duty, you've heard of it but probably don't realize it. It was a plane that was made of titanium. Titanium heats up really fast and expands so as it ascends into the sky (it can go pretty high, it's mach 3+; teacher said that the point where you can see stars in space during the day). But there was a technical problem because the heat was too much when it got up in the atmosphere so they had to use the fuel lines lining the walls of th titanium to cool it so that the titanium didn't begin to bend due to thermal expansion. here's the wikipedia page for it http://en.wikipedia.org/wiki/Lockheed_SR-71_Blackbird#SR-71

OpenStudy (anonymous):

expands so as it ascends into the sky , the wings actually get longer by a lot!

OpenStudy (anonymous):

but there was a problem with the lining because as the lining got hotter it also expanded so when the blackbird was low in the sky it actually leaked gas until it got higher into the atmosphere

OpenStudy (anonymous):

leaked gas due to the fact they had to make the wall linings small so it could expand

OpenStudy (anonymous):

ooooo... interesting (: I'm guessing your a gamer haha?!

OpenStudy (anonymous):

considering your major as well lol

OpenStudy (anonymous):

haha a bit .... i just remembered that when i used to play call of duty when i was younger , they had the blackbirds =D .. they technically don't make them anymore

OpenStudy (anonymous):

That's really interesting though (: Imma read more into that..... [after I study, of course] haha

OpenStudy (anonymous):

haha alright have fun studying =]

OpenStudy (anonymous):

thanks for your help!